# Tracking Any Object Amodally

### v2

@@ Abstract
Amodal perception, the ability to comprehend complete object structures from partial visibility, is a fundamental skill, even for infants. Its significance extends to applications like autonomous driving, where a clear understanding of heavily occluded objects is essential. However, modern detection and tracking algorithms often overlook this critical capability, perhaps due to the prevalence of modal annotations in most datasets. To address the scarcity of amodal data, we introduce the TAO-Amodal benchmark, featuring 880 diverse categories in thousands of video sequences. Our dataset includes amodal and modal bounding boxes for visible and occluded objects, including objects that are partially out-of-frame. To enhance amodal tracking with object permanence, we leverage a lightweight plug-in module, the amodal expander, to transform standard, modal trackers into amodal ones through fine-tuning on a few hundred video sequences with data augmentation. We achieve a 3.3% and 1.6% improvement on the detection and tracking of occluded objects on TAO-Amodal. When evaluated on people, our method produces dramatic improvements of 2x compared to state-of-the-art modal baselines.
@@ 1 Introduction
To address this gap, we first introduce a benchmark for large-scale amodal tracking, which requires estimating the full extent of objects through heavy and even complete occlusions. Our benchmark annotates 17,000 objects with amodal bounding boxes, along with human confidence estimates, from 880 classes in 2,921 videos. While prior datasets focus on images or are limited to a small vocabulary of classes (Tab. 1), our benchmark evaluates amodal tracking for hundreds of object classes. Since objects can get occluded because of other objects in the scene, and because of the limited field-of-view of cameras in casual captures, we define and address two kinds of occlusions – in-frame, and out-of-frame. As annotating amodal bounding boxes can be ambiguous and challenging, we design a new annotation protocol with detailed guidelines to improve human annotation. Importantly, we base our benchmark on a large-vocabulary multi-object tracking dataset, TAO [10]. This choice allows us to pair our amodal box annotations with class labels, modal boxes, and precise modal mask annotations [2] collected in prior work.
@@
Equipped with this data, we set out to evaluate the difficulty of amodal tracking. We evaluate using standard metrics, including detection and tracking AP, and variants that evaluate tracking specifically under partial and complete occlusions. As expected, we find that standard trackers trained with modal annotations do not suffice for amodal tracking.
We present an effective module to turn modal trackers into amodal trackers with object permanence. Specifically, we build a class-agnostic amodal expander, inspired by [65], that expands modal bounding box predictions to cover the full extent of objects. With limited amodal training data, we build the amodal expander to be class-agnostic, allowing us to transfer across classes for which we have fewer amodal annotations. Specifically, for amodal detection of people, we observe dramatic relative improvements of $\sim$30.0 AP points as compared to modal baselines. Our results show strong improvements over state-of-the-art trackers for amodal tracking.
@@
In summary, our contributions are as follows: (1) We annotate a large-scale dataset of amodal tracks for diverse objects, consisting of 17k objects spanning 880 categories, (2) we adapt evaluation metrics to handle amodal settings, and evaluate state-of-the-art trackers for our new task, and finally, (3) we present a strong baseline for amodal tracking with object permanence, using a category-agnostic amodal expander module.
# Sequences # Occluded Boxes # Occluded Track Ann Total
@@
Train Val Test Classes Partial Heavy OoF tracks length fps length
COCO-Amodal [65] 2500 1250 1250 $-*{}^{*}$ 8.8k 0.2k 0 - - - -
Sail-VOS [26] 160 41 0 162 559.5k 704.8k 0 7.9k 14.14 8 3,359
Sail-VOS-3D [27] 161 41 0 24 295.0k 387.5k 0 5.0k 13.10 8 2,808
NuScenes [5] 700 150 150 23 571.1k 139.5k 219k 24.5k 9.06 20 6,000
MOT17 [42] 7 0 7 1 51.2k 16.4k 16k 0.1k 6.98 30 248
MOT20 [11] 4 0 4 1 729.4k 88.1k 88k 1.6k 20.55 25 178
@@
500 993 1428 880 158.2k 35.1k 139k 9.6k 22.24 1 88,605
Table 1: Statistics of amodal datasets. We compare TAO-Amodal to prior image (first block), synthetic video (second block), and real video (last block) datasets. Compared to existing amodal datsets, TAO-Amodal is notable for being real-world videos that span far more categories and far more annotated frames for evaluation. We define heavy occlusion as objects with visibility below 10%, and partial as between 10%-80%. Occluded tracks are those that have heavy or partial occlusions for more than 5 seconds. Out-of-frame (OoF) objects are ones that extend partially beyond the image boundary. Track length is averaged over the dataset in seconds, while total length is the length of eval sequences in seconds.
@@ 2 Related work
Amodal perception has been studied in the past by benchmarks and algorithms, in both the single-frame (detection) and multi-frame (detection and tracking) settings. Since amodal object annotations are hard to obtain due to the uncertainty in human annotations (c.f. prior work [30] on a human vision experiment), the community has seen more dependence either on synthetic datasets, or real-world datasets with few classes and limited diversity. The remainder of the sections talk about this in detail.
@@ 2.1 Benchmarks
Real-world datasets
In the surveillance and self-driving domains, real-world datasets that annotate objects amodally exist. MOT 15-20 [33, 42, 11] evaluate multi-object tracking on amodal person detections obtained from detectors trained on MOT annotations. One pitfall for MOT is that instead of humans labelling the boxes during “invisible” phases of people, these box annotations are just linearly interpolated between two visible frames of the same object. Additionally, the metrics used by MOT weigh all modal and amodal annotations equally. This downweighs the tracking performance on amodal objects as they form only a small fraction of all annotations.
A number of multimodal (images and 3D LiDAR) datasets for autonomous driving have recently become popular. These include ArgoVerse (1.0 and 2.0) [7, 58], Waymo [50], nuScenes [5] and KITTI [18]. These datasets aim to focus on 3D tasks, and therefore use human annotators to label all objects in 3D to their full extent. In this setting, amodal perception is organic, and almost trivial, as in a fully-3D world there are no occlusions. These 3D boxes when projected onto 2D images would be useful for amodal perception but unfortunately, they only cover a small number of object classes. Another way to obtain amodal object annotations is in a multi-view setting. Datasets like CarFusion [45] and MMPTrack [22] follow this data curation scheme but because of the cumbersome data collection process, they are limited to only a single or few categories.
In the single-frame setting, COCO-Amodal, Amodal KINS and NuImages [5, 65, 44] contain amodal annotations of a few objects, but only cover the cases of partial occlusion (as complete occlusion is hard to resolve in a single-frame setting without temporal visibility cues). Moreover, COCO-Amodal [65] and Amodal KINS [44] are not shipped with class labels, which makes it difficult to learn object-specific priors for amodal completion.
Synthetic datasets
Intuitively, it is cheap to get amodal object annotations from synthetic data setups. SAIL-VOS and SAILVOS-3D [26, 27] are such datasets that exploit synthetic dataset curation and come with a number of different types of annotations (bounding boxes, object masks, object categories, their long-range tracks, and 3D meshes). Some of these even suit our case of detecting ‘out-of-frame’ occlusions, where one could project 3D meshes onto the image plane. While the number of categories are slightly larger for these datasets (including others like ParallelDomain [53] and DYCE [14]), the sim-to-real transfer remains a challenge even for modal perception [8, 29].
@@ 2.2 Algorithms
Based off of some amodal datasets, there has been a growing interesting in developing algorithms suitable for amodal perception. Some methods aim to track objects with object permanence [30, 53, 54, 55]. Some approaches utilize prior-frame information [63, 6, 60, 49, 13, 4, 59]. For instance, GTR [53] employs a transformer-based architecture and uses trajectory queries to group bounding boxes into trajectories. Previous work also segment objects amodally [35, 39]. We lean on similar approaches in this work, and devise a mechanism to generate occlusion cases in the flavor of the data augmentation used by GTR [64], and show that this is essential to the goal of enabling amodal perception.
@@ 3 Dataset Annotation
Existing datasets for modal perception are limited either in terms of their diversity, or the vocabulary of classes labelled. To this end, we build upon the modally annotated TAO dataset. It contains bounding box track annotations of 833 object categories at 1FPS spanning a total of 2,921 videos from 7 different data sources (AVA [20], Argoverse [7], Charades [47], HACS [62], LaSOT [16], BDD100K [61], YFCC100M [52]). This gives us the advantage of adding amodal box annotations to an already existing set of multimodal annotations in TAO (i.e., object classes, modal bounding boxes and modal segmentation masks). In terms of the extent/scope of annotation, TAO follows the single-frame detection datasets like LVIS and OpenImages [21, 31], in adopting a federated annotation protocol for object tracking, i.e., not every object is exhaustively annotated in every video. For collecting amodal annotations in TAO-Amodal, we consider only the set of tracks that are labelled in TAO.
Scope.
@@
Since annotators can exhibit a large variation in annotating the precise shape of objects while they undergo partial or even complete occlusion, we annotate using bounding boxes (instead of segmentation masks like in BURST [2]) to annotate their extent in the visible scene. Additionally, we define ‘in-frame’ occlusions as those occurring from the presence of occluders (occluders might be other dynamic objects, or static scene elements), and ‘out-of-frame’ occlusions as those resulting from objects leaving the camera field-of-view. We do not label the extent of occlusion in cases where an object may be partially present behind the camera (e.g., a person holding the camera who has their hands visible in the image). For labelling ‘out-of-frame’ occlusions, we need to fix bounds for annotation on the image plane. We ask annotators to work within an annotation workspace that extends to twice the image dimensions in consideration, with the image itself horizontally and vertically center-aligned in this workspace.
Figure 2: ROI Head [19] with Amodal Expander. Our Amodal Expander serves as a plug-in module that can “amodalize” any existing detector or tracker with limited (amodal) training data. It operates by taking as input region proposal features and modal box predictions (often represented as a residual delta with respect the region proposal) and generates amodal box outputs (again represented as residual deltas). To enable learning from limited data, we freeze all modules except the amodal expander.
Annotation Protocol
@@
Since TAO tracks (and their individual bounding boxes) are modal in nature, extending boxes to account for in-frame and out-of-frame occlusions translates to (1) modifying existing boxes in TAO, for the cases of partial occlusion, and (2) adding new boxes for completely occluded objects. We find that out of a total of 358,862 boxes in TAO, 266,902 (74.4%) boxes are modified to incorporate partial occlusion. TAO-Amodal boasts of an additional 23,449 bounding boxes for invisible objects that were not in TAO. These annotations follow the guidelines detailed in the appendix, that cover a wide range of both in-frame and out-of-frame occlusion scenarios. Importantly on a higher-level, we only consider the occlusion cases, where an object has appeared in the scene before. We also scope out the occlusion cases, where an object might be partially behind the camera, or outside the annotation workspace defined above. Within the strict purview of the guidelines, when an object’s location cannot be discerned confidently by the annotators, an is_uncertain flag is added to the labeled bounding box. From the 23,449 boxes for invisible objects, 20,218 (85.8%) boxes are annotated confidently (i.e., without the uncertain flag), indicating that there is inherent uncertainty in localizing objects when they undergo heavy occlusions (ref. human vision experiment from a prior work [30]).
@@
Finally, equipped with both modal and amodal annotations for all objects in TAO and TAO-Amodal, we add a visibility field to the TAO-Amodal annotations. From qualitative visual inspection, we find that the IoU between the modal and amodal boxes, is a good enough approximation of the visibility of each object.
Quality Control
@@
We ensure two rounds of professional quality check on TAO-Amodal annotations; this means all bounding box annotations are refined twice by annotators. Finally, authors conduct a manual quality check where 7 videos are chosen at random from the 7 dataset subsets and only 2 out of 349 tracks are found to be erroneous. These are tracks of invisible objects (visibility 0.0%) where some boxes are without a uncertainty flag. Our analysis show that nearly all inspected tracks ($>99\%$) are accurate, indicating the high-quality of amodal tracking annotations in TAO-amodal.
3.1 Statistical Analysis

### v3

# https://tao-amodal.github.io

@@ Abstract
Amodal perception, the ability to comprehend complete object structures from partial visibility, is a fundamental skill, even for infants. Its significance extends to applications like autonomous driving, where a clear understanding of heavily occluded objects is essential. However, modern detection and tracking algorithms often overlook this critical capability, perhaps due to the prevalence of modal annotations in most benchmarks. To address the scarcity of amodal benchmarks, we introduce TAO-Amodal, featuring 833 diverse categories in thousands of video sequences. Our dataset includes amodal and modal bounding boxes for visible and partially or fully occluded objects, including those that are partially out of the camera frame. We investigate the current lay of the land in both amodal tracking and detection by benchmarking state-of-the-art modal trackers and amodal segmentation methods. We find that existing methods, even when adapted for amodal tracking, struggle to detect and track objects under heavy occlusion. To mitigate this, we explore simple finetuning schemes that can increase the amodal tracking and detection metrics of occluded objects by 2.1% and 3.3%.
Keywords:
Amodal perception Large-scale evaluation benchmark Multi-object tracking.
Figure 1: TAO-Amodal. We present TAO-Amodal, a dataset of amodal (bounding box) annotations for fully occluded and partially occluded (both within the image frame and out-of-frame) objects in videos from the TAO dataset [10]. Our dataset consists of 332k boxes that cover multiple occlusion scenarios across 2,907 videos with annotations for 833 object categories. TAO-Amodal aims at assessing the occlusion reasoning capabilities of current trackers for amodal tracking of any object.
@@ 1 Introduction
To address this gap, we introduce a benchmark for large-scale amodal tracking, which requires estimating the full extent of objects through heavy and even complete occlusions. Our benchmark annotates 17,000 objects with amodal bounding boxes, along with human confidence estimates, from 833 classes in 2,907 videos. While prior datasets focus on images or are limited to a small vocabulary of classes (Tab. 1), our benchmark evaluates amodal tracking for hundreds of object classes. Since objects can get occluded because of other objects in the scene, and because of the limited field-of-view of cameras in casual captures, we define and address two kinds of occlusions – in-frame, and out-of-frame. As annotating amodal bounding boxes can be ambiguous and challenging, we design a new annotation protocol with detailed guidelines to improve human annotation. Importantly, we base our benchmark on a large-vocabulary multi-object tracking dataset, TAO [10]. This choice allows us to pair our amodal box annotations with class labels, modal boxes, and precise modal mask annotations [2] collected in prior work.
@@
Equipped with this data, we set out to evaluate the difficulty of amodal tracking. We evaluate using standard metrics, including detection and tracking AP, and variants [33] that evaluate tracking specifically under partial and complete occlusions. As expected, we find that standard trackers trained with modal annotations do not suffice for amodal tracking.
To adapt existing modal trackers into amodal ones, we finetune them on TAO-Amodal. The closest line of work to amodal tracking is amodal segmentation [66, 38, 47]. We benchmark recent amodal segmentation algorithms by running a Kalman-Filter based association during post-processing on their predictions. While this addresses the gap between modal and amodal tracking to some extent, the performance is far from good due to the challenging occlusion scenarios in TAO-Amodal. To mitigate this, we explore different but simple finetuning and data-augmentation strategies inspired by prior work [38, 72]. This lets us set a new baseline on the tasks of amodal detection and tracking.
@@
In summary, our contributions are as follows: (1) we annotate a large-scale dataset of amodal tracks for diverse objects, consisting of 17k objects spanning 833 categories, (2) we adapt evaluation metrics to handle amodal settings, and evaluate state-of-the-art trackers for our new task, and finally, (3) we investigate multiple finetuning and data-augmentation schemes as simple extensions to improve the existing modal tracking algorithms.
Table 1: Statistics of amodal datasets. TAO-Amodal is proposed as an evaluation benchmark for amodal tracking. We compare our dataset to prior image (first block), synthetic video (second block), and real video (last block) datasets. TAO-Amodal is notable for being real-world videos that span far more categories and far more annotated frames for evaluation. Track length is averaged over the dataset in seconds, while total length is the length of eval sequences in seconds. We define heavy occlusion as objects with visibility below 10%, and partial as between 10%-80%. Occluded tracks are those that have heavy or partial occlusions for more than 5 seconds. Out-of-frame (OoF) objects are ones that extend partially beyond the image boundary.
# Sequences Track Total # Occluded Boxes # Occluded Ann
@@
Total Test Val Train Classes length length Partial Heavy OoF tracks fps
COCO-Amodal [72] 5000 1250 1250 2500 - - - 8.8k 0.2k 0 - -
Sail-VOS [29] 201 0 41 160 162 14.14 3,359 559.5k 704.8k 0 7.9k 8
Sail-VOS-3D [30] 202 0 41 161 24 13.10 2,808 295.0k 387.5k 0 5.0k 8
NuScenes [5] 1000 150 150 700 23 9.06 6,000 571.1k 139.5k 219k 24.5k 20
MOT17 [45] 14 7 0 7 1 6.98 248 51.2k 16.4k 16k 0.1k 30
MOT20 [11] 8 4 0 4 1 20.55 178 729.4k 88.1k 88k 1.6k 25
@@
2907 1419 988 500 833 22.24 88,605 158.2k 35.1k 139k 9.6k 1
Figure 3: Class distribution. We present counts of instances from top 8 most frequent categories and other categories, using a logarithmic scale.
Figure 4: Object occlusion distribution. We plot the distribution at a 10% visibility span.
@@ 2 Related work
Amodal perception has been studied in the past by benchmarks and algorithms, in both the single-frame (detection) and multi-frame (detection and tracking) settings. Since amodal object annotations are hard to obtain due to the uncertainty in human annotations (c.f. prior work [33] on a human vision experiment), the community has depended heavily on synthetic datasets, or real-world datasets with few classes and limited diversity. We provide an overview of this prior wrok in the rest of this section.
@@ 2.1 Benchmarks
2.1.1 Real-world datasets.
Amodal object annotations for real-world scenes are largely limited to the surveillance and self-driving domains.MOT 15-20 [36, 45, 11] evaluate multi-object tracking on amodal person detections obtained from detectors trained on MOT annotations. However, these amodal annotations are automatically propagated via linear interpolation of annotations in frames where objects are visible. Additionally, the metrics used by MOT weigh all modal and amodal annotations equally. This largely ignores tracking performance on amodal objects, which form only a small fraction of all annotations.
A number of multimodal (images and 3D LiDAR) datasets for autonomous driving have recently become popular. These include ArgoVerse (1.0 and 2.0) [7, 62], Waymo [53], nuScenes [5] and KITTI [19]. These datasets aim to focus on 3D tasks, and therefore use human annotators to label all objects in 3D to their full extent. In this setting, amodal annotations arise naturally due to the 3D nature of the data. These 3D boxes, when projected onto 2D images, would be useful for amodal perception; unfortunately, these annotations cover only a small number of object classes. Another way to obtain amodal object annotations is in a multi-view setting. Datasets like CarFusion [48] and MMPTrack [24] follow this data curation scheme, but, due to the cumbersome data collection process, they are limited to only a single or few categories.
In the single-frame setting, COCO-Amodal, Amodal KINS and NuImages [5, 72, 47] contain amodal annotations, but only cover the cases of partial occlusion: complete occlusions can only be recovered with temporal information, which is missing in image datasets. Moreover, COCO-Amodal [72] and KINS [47] do not provide class labels, which makes it difficult to learn object priors for amodal completion, and to evaluate the accuracy of amodal tracking in the long tail.
Synthetic datasets. An alternative approach to the above is use synthetic data generation pipelines to get amodal annotations. SAIL-VOS and SAILVOS-3D [29, 30] are such datasets that exploit synthetic dataset curation and come with a number of different types of annotations (bounding boxes, object masks, object categories, their long-range tracks, and 3D meshes). Some of these even suit our case of detecting ‘out-of-frame’ occlusions, where one could project 3D meshes onto the image plane. While the number of categories are slightly larger for these datasets (including others like ParallelDomain [57] and DYCE [14]), the sim-to-real transfer remains a challenge even for modal perception [8, 32].
@@ 2.2 Algorithms
Based off of some amodal datasets, there has been a growing interesting in developing algorithms suitable for amodal perception. Some methods aim to track objects with object permanence [33, 57, 56, 59]. Previous work also segment objects amodally [38, 42, 66, 67, 18]. Some approaches utilize prior-frame information [70, 6, 64, 52, 13, 4, 63, 71]. For instance, GTR [71] employs a transformer-based architecture and uses trajectory queries to group bounding boxes into trajectories. We lean on similar approaches in this work, and devise a mechanism to generate occlusion cases in the flavor of the data augmentation used by GTR [71], and show that this is essential to the goal of enabling amodal perception.
Figure 5: ROI Head [21] with Amodal Expander. Amodal Expander serves as a plug-in fine-tuning scheme to “amodalize" existing detectors or trackers with limited (amodal) training data. It operates by taking as input region proposal features and modal box predictions (often represented as a residual delta with respect the region proposal) and generates amodal box outputs (again represented as residual deltas). We freeze all modules except the amodal expander during fine-tuning.
Figure 6: Amodal Expander Architecture. Given $N$ flattened proposal features and modal box (delta) predictions represented with 256-dim positional encodings [60], we predict amodal box (deltas) with a two-layer MLP (unless otherwise specified). Further architecture details are in Sec. 5.2.
Table 2: Amodal trackers on TAO-Amodal validation set. We define metrics in Sec. 5.1. The visibility range is indicated by the superscript to denote various levels of occlusion. We fine-tuned modal trackers on TAO-Amodal-train for 20k iterations. Detector [40] and amodal segmentation methods [67, 18, 58] were evaluated using Kalman filter based association [4]. We evaluated models that predict COCO vocabulary [41] using objects within COCO category. GTR is used as the basis for subsequent experiments, considering its performance in detection and tracking metrics. We run all trackers at 1 fps and average AP across categories with an IoU threshold of 0.5.
Detection Metrics Tracking Metrics
[…]
PCNet [67] 0.48 7.15 21.43 8.69 15.59 5.80 3.91
QDTrack [17] 0.35 8.03 21.82 8.05 15.62 7.84 4.03
TET [39] 0.24 5.77 14.98 4.87 10.86 4.84 3.44
ViTDet-B [40] 0.77 12.57 34.33 14.18 25.94 7.66 4.38
ViTDet-L [40] 1.25 15.06 38.16 15.84 29.04 9.70 5.90
ViTDet-H [40] 1.13 15.80 40.09 16.97 30.20 9.72 5.63
GTR [71] 0.77 14.62 38.17 15.31 29.24 16.07 9.28
COCO category eval
ORCNN [18] 0.33 11.78 37.88 16.68 26.09 5.72 3.43
AmodalMRCNN [18] 0.46 14.74 42.65 18.35 29.58 7.57 4.47
AISFormer [58] 0.36 14.23 39.76 18.61 27.70 7.88 5.90
PCNet [67] 1.30 20.61 53.13 24.21 37.04 11.19 8.78
Table 3: Exploring fine-tuning strategies on TAO-Amodal validation set. We ablate different strategies for repurposing GTR for amodal tracking. Fine-tuning plug-in expander modestly outperforms fine-tuning all or part of the model. Combined with data augmentation, PasteNOcclude (PnO), for generating synthetic occlusions (as detailed in Sec 4.2), expander produces noticeable gains for partially occluded and out-of-frame objects. All models (other than the baseline) were trained on TAO-Amodal training set for 20k iterations, while denotes 45k iterations of training.
Detection Metrics Tracking Metrics
[…]
Baseline (GTR [71]) 0.78 13.24 37.54 14.18 28.19 16.02 8.86
Fine-tune entire model 0.52 10.36 24.08 10.34 17.93 7.70 3.93
FT entire model + PnO 0.79 9.68 26.56 10.10 20.16 9.05 4.30
Fine-tune regression head
& proposal network
0.79 10.57 27.91 11.37 21.42 9.04 4.53
Fine-tune regression head 0.77 14.62 38.17 15.31 29.24 16.07 9.28
FT regression + PnO 0.87 14.36 38.18 15.47 29.04 15.95 9.23
Amodal Expander 0.67 16.29 37.11 17.39 29.50 16.10 10.44 (+1.58)
Amodal Expander + PnO 0.77 16.53 (+3.29) 37.80 (+0.26) 17.65 (+3.47) 29.96 (+1.77) 16.35 (+0.33) 10.28 (+1.42)
@@ 3 Dataset Annotation
Existing datasets for modal perception are limited either in terms of their diversity, or the vocabulary of classes labelled. To this end, we build upon the modally annotated TAO dataset. It contains bounding box track annotations of 833 object categories at 1FPS spanning a total of 2,921 videos from 7 different data sources (AVA [22], Argoverse [7], Charades [50], HACS [68], LaSOT [16], BDD100K [65], YFCC100M [55]). Bootstrapping from this dataset allows us to add amodal box annotations to an already existing set of multimodal annotations in TAO – i.e., object classes, modal bounding boxes and modal segmentation masks. TAO follows the single-frame detection datasets, such as LVIS and OpenImages [23, 34], in adopting a federated annotation protocol for object tracking: i.e., not every object class is exhaustively annotated in every video. We refer the reader to [10, 23] for details on federated annotation and evaluation setup, and focus here on our amodal annotation of objects in TAO.
@@
Scope. Since annotators can exhibit a large variation in annotating the precise shape of objects while they undergo partial or even complete occlusion, we annotate using bounding boxes instead of segmentation masks to mark the full extent of objects in the visible scene. We define ‘in-frame’ occlusions as those occurring from the presence of occluders (which may be other dynamic objects, or static scene elements), and ‘out-of-frame’ occlusions as those resulting from objects leaving the camera field-of-view. We do not label the extent of occlusion in cases where an object may be partially present behind the camera (e.g., a person holding the camera who has their hands visible in the image). For labelling ‘out-of-frame’ occlusions, we need to fix bounds for annotation on the image plane. We ask annotators to work within an annotation workspace that extends to twice the image dimensions in consideration, with the image itself horizontally and vertically center-aligned in this workspace.
@@
Annotation Protocol. Since object tracks in TAO are modal in nature, extending boxes to account for in-frame and out-of-frame occlusions requires (1) (in the case of partial occlusion) complementing TAO bounding boxes with amodal boxes, and (2) (in the case of complete occlusion) adding new boxes to object tracks for occluded frames. Out of a total of 358,862 boxes in TAO, our annotators modify 266,902 (74.4%) to account for partial occlusions. Further, TAO-Amodal introduces an additional 23,449 bounding boxes for frames where objects were invisible and unlabeled in TAO. These annotations follow the guidelines detailed in the appendix, covering a wide range of both in-frame and out-of-frame occlusion scenarios. Importantly, we only consider occlusion cases where an object has appeared in the scene before. We exclude occlusions where an object might be partially behind the camera or outside the annotation workspace defined above. Within the strict purview of the guidelines, when an object’s location cannot be discerned confidently by the annotators, annotators are instructed to mark an is_uncertain flag. From the 23,449 boxes for invisible objects, 20,218 (85.8%) boxes are annotated confidently (i.e., without the uncertain flag), indicating that there is inherent uncertainty in localizing objects when they undergo heavy occlusions (matching observations from prior work [33] which indicate human uncertainty of object location under occlusion).
@@
Finally, equipped with both modal and amodal annotations for all objects, we add a visibility field to the TAO-Amodal annotations, using the overlap (intersection-over-union) between the modal and amodal boxes as a proxy.
@@
Quality Control. We conduct two rounds of professional quality checks on TAO-Amodal annotations: all bounding box annotations are refined twice by annotators. Finally, the authors of this work conducted a manual quality check reviewing 349 tracks from 7 randomly sampled videos, and found only 2 (<1%) tracks without an uncertainty flag to be erroneous. Both tracks were for objects with complete occlusions (visibility 0.0%) in the video. Our analysis show that nearly all inspected tracks ($>99\%$) are accurate, indicating the high-quality of amodal tracking annotations in TAO-amodal.
3.1 Dataset statistics